OpenStudy (djbreezy):

Which pairs are solutions to the equation? 5xy+9=44 A. (1,7) and (4,11) B. (1,7) and (7,1) C. (7,1) and (2,3) D. (4,11) and (2,3)

OpenStudy (djbreezy):

@ILovePuppiesLol

ILovePuppiesLol (ilovepuppieslol):

again we go by trying out the options i will try option B ;) \(5xy+9=44\) put (1,7) \(5 \times 1 \times 7 +9=44\) \(35=44\) \(44=44\) true put (7,1) \(5 \times 7 \times 1 +9=44\) \(35=44\) \(44=44\) true so this si correct

OpenStudy (djbreezy):

Do u mean B??

ILovePuppiesLol (ilovepuppieslol):

yee